Sove for X. write in decimal notation. (73/24)^x=4/135
Answer by edjones(8007)

(73/24)^x=4/135
log[73/24]^x=log[73/24](4/135)
x=log[73/24](4/135)
log[73/24](4/135)=log[10](4/135)/log[10](73/24) change of base formula.
x=-3.1634
.
Ed